### v1

@@ 3.3 Related Comparison
Attribute
SPINEPS (ours)
SPIDER nnUNet baseline (18)
Li et al. (11)
Sáenz-Gamboa et. al. (9)
Input MRI
T2w sagittal
T2w sagittal
T2w axial
T1w and T2w sagittal (requires both)
Dimensionality
3D
3D
2D slices
2D slices
Outputs semantic mask
Yes
No
Yes
Yes
Outputs instance mask
Yes
Yes
No
No
Lumbar region
Yes
Yes
Yes
Yes
Thoracic region
Yes
No
No
No
Cervical region
Yes
No
No
No
Vertebra Corpus DSC ↑
0.94
N/A
0.98
N/A
Vertebra Corpus ASSD ↓
0.34
N/A
2.32
N/A
Global Vertebra DSC ↑
0.92
0.92
N/A
0.93
Global IVD DSC ↑
0.96
0.89
N/A
0.94
Global spinal canal DSC ↑
0.95
0.92
N/A
0.88
Includes posterior elements
Yes
Yes
No
Yes
Includes vertebra substructures
Yes
No
No
No
Publicly available
Yes
Yes
No
No

### v2

@@ 3.3 Related Comparison
Attribute SPINEPS (ours) SPIDER nnUNet baseline (18) Li et al. (11) Sáenz-Gamboa et. al. (9)
Input MRI T2w sagittal T2w sagittal T2w axial T1w and T2w sagittal (requires both)
Dimensionality 3D 3D 2D slices 2D slices
Outputs semantic mask Yes No Yes Yes
Outputs instance mask Yes Yes No No
Lumbar region Yes Yes Yes Yes
Thoracic region Yes No No No
Cervical region Yes No No No
Vertebra Corpus DSC ↑ 0.94 N/A 0.98 N/A
Vertebra Corpus ASSD ↓ 0.34 N/A 2.32 N/A
Global Vertebra DSC ↑ 0.92 0.92 N/A 0.93
Global IVD DSC ↑ 0.96 0.89 N/A 0.94
Global spinal canal DSC ↑ 0.95 0.92 N/A 0.88
Includes posterior elements Yes Yes No Yes
Includes vertebra substructures Yes No No No
Publicly available Yes Yes No No